# Physics-Informed Weakly Supervised Learning for Interatomic Potentials

### v1

@@ 5.3 Qualitative Impact of PIWSL
We evaluate the prediction variance and robustness of an MLIP model trained with PIWSL using the aspirin molecule, focusing on the potential energy’s dependence on the C–H bond length. In this work, robustness refers to the prediction robustness of an MLIP to perturbations in atomic coordinates. In literature, the robustness of MLIPs often means their stability during MD simulations, which is outside the scope of this paper. We train PaiNN on the rMD17 aspirin data set using $100$ and $200$ configurations with and without the PIWSL loss. The detailed training setup and errors of the used MLIPs are summarized in Section D.4. We examine the potential energy varying the length of a C–H bond from $0.9$ to $1.4$. The equilibrium C–H bond length is about $1.09$. The results in Fig. 2 (c) and (d) show that the PIWSL method improves the predicted potential energy profile, demonstrating increased robustness under perturbations of atom coordinates. This improvement in the potential curve prediction is achieved without introducing unphysical noise, which can occur when using inappropriate weak labels for the perturbed atomic coordinates.
@@
Although the estimated potential energies do not always match the reference values, the direction between the original and perturbed configurations, indicated by arrows in Fig. 2, consistently follows the gradient of the reference potential energy, corresponding to the negative force. In Fig. 2, we use a perturbation length of $||\delta\mathbf{r}||=0.01$. This consistency with the energy gradient underscores the PIWSL method’s effectiveness, ensuring the alignment between predicted total energies and force and improving the corresponding RMSE values. As discussed in Section 4.2, PIWSL also addresses the limitation of MLIPs that employ separate force branches and do not guarantee the prediction of conservative forces. The proposed method reduces the curl of predicted forces, as detailed in Section D.11, although complete elimination of the curl remains a challenge. In summary, PIWSL minimizes individual energy and force errors, improving the overall accuracy of MLIPs.
5.4 Training MLIPs without Reference Forces
In the following, we explore scenarios where only potential energy labels are available. This situation commonly arises when calculating energy labels with chemically accurate approaches, such as CCSD(T)/CBS [21, 22], for which force calculation is infeasible. To consider practical applications, we examine two cases: (1) predicting force by a force branch (FB) and (2) predicting force as a gradient of the potential energy (GF). The former enables fast force prediction and is popular in the machine learning community, while the latter requires additional gradient calculation but yields curl-free force predictions. It is popular in computational chemistry as it ensures the conservation of the total energy during MD simulations. The results are provided in Table 3; training without reference forces is achieved by setting the relative force contribution to zero in Eq. (1). The PIWSL method consistently performs better than the baseline for the FB and GF cases. However, a more significant improvement in the force prediction performance is observed in the GF case. We attribute this phenomenon to the inherent nature of PIWSL, which requires consistency between the potential energy and atomic forces, as discussed in Section 5.3. This result aligns with our expectations, confirming the capability of our PIWSL method to enable ML models to reduce the error in the predicted forces. Overall, PIWSL opens a new possibility for training MLIP models using highly accurate reference methods, such as CCSD(T)/CBS.
5.5 Further Analyses of PIWSL
Comparing PITC with the Taylor-Expansion-Based Weak Label Loss. We compare our proposed method with the Taylor-expansion-based weak label (WL) approach [17]. For the WL method, we utilize Eq. (A3) as the loss function. For simplicity, we only consider the PITC loss in Eq. (3). For a fair comparison, we consider the following two cases. First, we train with reference forces and energies (w. RF). Second, we train the methods without reference forces and use only the reference energies. For the training with reference forces, we set the numeric coefficient of the PITC loss to $1.0$; for the training without reference forces [^1] Calculating the WL loss without the reference force is impossible when force labels are unavailable [17]. The additional experiment of the WL loss with the force prediction is presented in Section D.10., the coefficient is set to $0.1$. The results are provided in Table 5.
Our PITC loss demonstrates the best accuracy in all cases, with and without the reference forces. Interestingly, PaiNN failed to learn the potential energy with the WL loss and reference forces. We hypothesize this to be due to the imbalance of the training between the energies and forces. Specifically, the WL loss trains only the potential energy, resulting in an inconsistency between the energy and force branches, which share the same readout layer that experiences more frequent updates using the potential energy. This hypothesis is supported by the results for the training without reference forces, where the error in energy is reduced compared to the baseline. A further validation in a similar experiment in the case of GF is provided in Section D.9. However, the proposed PITC loss still performs better here. In summary, the PITC loss enables MLIPs to learn energies and forces consistent with each other and does it better than the previously proposed WL method.
Table 5: Comparison of PITC and the Taylor-expansion-based weak label loss. WL denotes the Taylor-expansion-based method using reference energies and atomic forces; see Eq. (A3). The listed values are the RMSE values for energies in kcal/mol and atomic forces in kcal/mol/Å. All models are trained on the ANI-1x data set using 1000 configurations, with (w.) and without (w/o.) reference atomic forces (RF).
Model Case Baseline PITC WL
PaiNN (w. RF) E 56.62 $\pm$ 2.80 30.94$\pm$ 0.56 81.86$\pm$ 9.39
F 12.96$\pm$ 0.06 12.04$\pm$ 0.04 14.54$\pm$ 0.12
(w/o. RF) E 42.36$\pm$ 0.30 25.42$\pm$ 0.72 35.79$\pm$ 0.70
F 24.25$\pm$ 0.00 20.54$\pm$ 0.08 24.25$\pm$ 0.01
Equiformer (w. RF) E 54.52$\pm$ 4.52 23.16$\pm$ 0.19 31.02$\pm$ 3.99
F 10.10$\pm$ 0.00 10.03$\pm$ 0.05 13.43$\pm$ 0.92
(w/o. RF) E 43.14$\pm$ 0.86 29.48$\pm$ 0.51 30.08$\pm$ 1.97
F 24.25$\pm$ 0.00 21.99$\pm$ 0.49 24.25$\pm$ 0.00
Ablating the Impact of PITC and PISC Losses. We conduct an ablation experiment to analyze the impact of PITC and PISC losses. Results in Table 5.2 indicate that the PITC loss predominantly improves the accuracy of resulting models, especially for PaiNN. Using just the PISC loss does not consistently improve accuracy but stabilizes training when combined with PITC. This combined approach notably benefits Equiformer v2. For Equiformer v2, we repeated the experiment five times to reduce the effect from an outlier for the PITC loss.
Adversarial Directions for Perturbing Atomic Positions. The following discusses the dependence of the PIWSL’s performance on selecting the vector $\delta\mathbf{r}$ in Eq. (5) employed to perturb atomic positions. The detailed implementation and setups are provided in Section B.1. Table 6 compares the results obtained for a randomly-sampled vector $\delta\mathbf{r}$ and for the one determined adversarially. The results demonstrate that both approaches improve the performance compared to the baseline without weak supervision, though the results might depend on the employed model.
Table 6: Dependence of the PIWSL’s performance on the atomic position perturbation vector. The listed numerical values are the root mean square errors (RMSEs) for the energy in kcal/mol and force in kcal/mol/Å. All results are provided for the ANI-1x data set and models trained using 1000 configurations.
Baseline Random (Eq. (6)) Adversarial (Eq. (7))
PaiNN E 56.62 $\pm$ 2.80 24.53 $\pm$ 0.48 33.67 $\pm$ 1.12
F 12.96 $\pm$ 0.18 11.43 $\pm$ 0.05 12.74$\pm$ 0.14
Equiformer E 54.52 $\pm$ 4.52 23.16 $\pm$ 0.50 20.54$\pm$ 0.21
F 10.10 $\pm$ 0.00 10.03 $\pm$ 0.03 9.93$\pm$ 0.04
@@ 6 Discussion and Limitations
This work introduces the PIWSL method, encompassing two distinct physics-informed weakly supervised loss functions, for learning MLIPs. These losses provide the physics-informed weak labels based on the Taylor expansion (PITC loss) and the spatial-consistency (PISC loss) of the potential energy. These physics-informed weak labels enable any MLIP to improve its accuracy, particularly in scenarios characterized by small training data set sizes. Such scenarios are common when investigating a new molecular or material system. Our extensive experiments demonstrate notable efficacy and efficiency of our method from various aspects: (i) dependence on the training data set size, (ii) the potential energy prediction variance and robustness in terms of a perturbation on a C–H bond length, and (iii) selection of the perturbation vector. In particular, it is shown that our PIWSL method enables ML models to improve the force prediction even without force labels, thereby opening a new possibility for training MLIPs using highly accurate reference methods, such as CCSD(T)/CBS.
Limitations. The proposed PIWSL method is tailored to ML models that predict atomic forces and total energies of atomic systems. It cannot be applied to other ML problems unrelated to computational chemistry or materials science applications. Although this work uses the first-order Taylor expansion to obtain weak labels in Eq. (2), employing a more sophisticated higher-order ordinary differential equation solver is a viable alternative.
Code availability
The source code will be made publicly available upon publication.
@@ Acknowledgements
Appendix A Related Work
Addressing Data Sparsity in MLIPs. Generating training data sets suitable for learning reliable MLIPs is challenging, especially when considering unexplored molecular and material systems. Numerous computationally expensive calculations with either ab initio or first-principles approaches are required for the latter. To mitigate this challenge, active learning (AL) methods, which utilize prediction uncertainty, can be applied [14, 52, 15, 53, 54, 16]. Furthermore, equivariant MLIPs often reduce required training data set sizes through improved data efficiency [38, 40].
Appendix B Experimental Setup, Baselines, and Data Sets
B.1 Experimental Setup
Code for Experiments. The code used to run our experiments builds upon the recent work [55] and extends it to integrate the latest Open Catalyst Project code [9]. We adopt hyper-parameters from the Open Catalyst (OC) project, tuned to the corresponding OC data set. Note that we do not use this data set in the presented work, whose main focus is training general-purpose MLIPs that can be used to run molecular dynamics (MD) simulations and geometry optimization. However, the OC data set has been designed to investigate the latter, making it less suitable for the current study. Our modifications include adjusting the learning rate scheduler, details of which can be found in our repository. For potential energy and force prediction, we utilize mean-absolute error (MAE) and $L_{2}$-norm (L2MAE) losses with coefficients of 1 and 100, respectively. More details on the model hyperparameters are provided in our repository. For the PITC and PISC loss functions, we use the mean square error (MSE) loss based on an experiment in Section D.5.
@@ B.1 Experimental Setup
a As explained in Table 2, the mini-batch size of the SchNet model was changed to 32 due to high RMSE values observed with a mini-batch size of four as the training data set size increased. A more detailed discussion of the results for SchNet is provided in Section D.1.
@@
The remaining hyper-parameters are the coefficients for the PITC and PISC losses ($C_{\rm PITC},C_{\rm PISC}$) and the maximum magnitude $\epsilon_{\mathrm{max}}$ of the perturbation vector $\delta\mathbf{r}$; see Table A3. These hyper-parameters are tuned using Optuna [56] for PaiNN and Equiformer v2. We used 1000 configurations drawn randomly from the original ANI-1x data set for training. Due to multiple local minima, Optuna identified several optimal hyper-parameter sets in each run. We selected the following representative combinations ($C_{\rm PITC},C_{\rm PISC},\epsilon_{\mathrm{max}})$ = Case A: (1.2, 0.8, 0.025), Case B: (1.0, 0, 0.01), Case C: (0.1, 0.01, 0.01), Case D: (1.2, 0.01, 0.025), Case E: (1.2, 0.01, 0.01), Case F: (1.2, 0.01, 0.015), Case G: (0.01, 0.001, 0.025), and Case H: (0.1, 0.01, 0.025).
Table A3: Hyper-parameters for the PIWSL loss. We selected the following hyper-parameter combinations using Optuna [56]: ($C_{\rm PITC},C_{\rm PISC},\epsilon_{\mathrm{max}})$= Case A: (1.2, 0.8, 0.025), Case B: (1.0, 0, 0.01), Case C: (0.1, 0.01, 0.01), Case D: (1.2, 0.01, 0.025), Case E: (1.2, 0.01, 0.01), Case F: (1.2, 0.01, 0.015), Case G: (0.01, 0.001, 0.025), and Case H: (0.1, 0.01, 0.025).
@@
Splitting Data Sets. We split the original data sets into training, validation, and test sets for our experiments. We shuffled the original data sets using a random seed and selected the training data sets of predefined sizes. For validation, we selected the same number of configurations as in the training data set if it exceeded 100 configurations; otherwise, we used 100 configurations to ensure sufficient validation size. For the rMD17 data set, following [55], we used 9000 configurations as a validation data set and another 10,000 for testing. We used the same test data set across different sizes of the training data sets for a fair performance comparison. We used 10,000 test configurations for ANI-1x and 1000 for TiO2 and LMNTO.
Training with Adversarial Directions. In our experiments, which defined perturbation vectors adversarially (see Section 5.5), we determined adversarial directions using Eq. (7). More concretely, we only considered the potential energy, i.e., $\mathbf{y}_{\rm pred}$ and $\mathbf{y}_{\rm label}$, to avoid Hessian calculations. In addition, we considered the loss function for the potential energy as $L_{\rm dist}$ in Eq. (7). The expression of $\mathbf{g}=\nabla_{\mathbf{r}}L_{\rm dist}$ is then
@@
Note that we used the relation: $\nabla_{\mathbf{r}_{i}}E^{\rm ref}_{\mathcal{S}}=-\mathbf{F}^{\rm ref}_{i,\mathcal{S}}$, to obtain the final expression. Though $E^{\rm ref}$ can also be interpreted as a constant regarding atom positions. We have chosen this expression to avoid the case where the adversarial direction $\mathbf{g}$ points to $\mathbf{F}^{\rm ref}_{i,\mathcal{S}}$ other than the very beginning of the training. Our experiments indicate that the employed expression is slightly better than its alternative. In our experiment, we also randomly flip the sign of $\mathbf{g}_{\mathcal{S}}$ to avoid overfitting to adversarial directions.
B.2 Baseline Methods
Data Augmentation with NoisyNode. In our experiments, we used the NoisyNode approach [43] as one of the baseline methods. This method aims to improve the performance of ML models by adding a perturbation to node features, i.e., atomic coordinates, and makes ML models recover original labels. This approach enables ML models to be more robust to noise in the data. Although the original method recommends adding a decoder network to learn the denoising process, we do not utilize it following previous work [37] and add the perturbation vector to atomic coordinates similar to PIWSL losses, fixing energy and force labels. We implement the NoisyNode approach in our code. Thus, we can expect slightly different behavior compared to the recent work [43, 37].
Taylor-Expansion-Based Weak Labels. Recent work proposed a similar Taylor-expansion-based weak label approach [17]. Nonetheless, the loss is different from the one in Eq. (3) as the authors used reference energy and atomic force labels to estimate weak energy labels $E^{\mathrm{ref}}_{\mathcal{S}_{\delta\mathbf{r}}}$ for perturbed atomic configurations $\mathcal{S}_{\delta\mathbf{r}}$

### v2

@@ 5.3 Qualitative Impact of PIWSL
Figure 3: Stability analysis of the MLIP models during MD simulations. Stability during MD simulations is assessed for the baseline MLIP models and those trained with PIWSL. Left: Models with the direct force branch. Right: Models with forces computed as negative gradients of the energy. All results are obtained for the aspirin molecule and MD simulations in the microcanonical (NVE) statistical ensemble. We measure stability during MD simulations according to Fu et al. (2023).
@@
We evaluate the prediction variance and robustness of an MLIP model trained with PIWSL using the aspirin molecule, focusing on the potential energy’s dependence on the C–H bond length. In this work, robustness refers to the prediction robustness of an MLIP to perturbations in atomic coordinates. In the literature, the robustness of MLIPs also means their stability during MD simulations. We train PaiNN on the rMD17 aspirin data set using $100$ and $200$ configurations with and without the PIWSL loss. The detailed training setup and errors of the used MLIPs are summarized in section E.4.1. We examine the potential energy varying the length of a C–H bond from $0.9$ to $1.4$. The equilibrium C–H bond length is about $1.09$. The results in figure 2 (c) and (d) demonstrate that the PIWSL method improves the predicted potential energy profile, indicating improved robustness under perturbations of atom coordinates.
@@
Although the estimated potential energies do not always match the reference values, the direction between the original and perturbed configurations, indicated by arrows in figure 2, consistently follows the gradient of the reference potential energy, corresponding to the negative force. In figure 2, we use a perturbation length of $||\delta\mathbf{r}||=0.01$. This consistency with the energy gradient underscores the PIWSL method’s effectiveness, ensuring the alignment between predicted total energies and force and improving the corresponding RMSE values. As discussed in section 4.2, PIWSL also addresses the limitation of MLIPs that employ separate force branches and do not guarantee the prediction of conservative forces. The proposed method reduces the curl of predicted forces, as detailed in section E.11, although complete elimination of the curl remains a challenge. In summary, PIWSL minimizes individual energy and force errors, improving the overall accuracy of MLIPs.
To further assess PIWSL’s impact, we evaluate the robustness during MD simulations of the MLIP models trained with and without PIWSL. We consider MD simulations of the aspirin molecule, with corresponding results presented in figure 3. We measured stability following the approach proposed in Fu et al. (2023). A detailed experimental setup is provided in section E.4.2. The results demonstrate that PIWSL improves the stability of MD simulations for both the direct and gradient-based force prediction models. The simulation times in figure 3 are shorter than those reported by Fu et al. (2023). This difference arises from our choice to perform MD simulations in the microcanonical (NVE) statistical ensemble instead of the canonical (NVT) statistical ensemble in Fu et al. (2023) to assess stability more accurately without the influence of a thermostat. Results for MD simulations conducted in the canonical statistical ensemble are provided in section E.4.2.
5.4 Fine-Tuning of Foundation Models
In this section, we present a case study demonstrating the application of PIWSL to the fine-tuning of foundation models using a sparse data set. The detailed experimental setup is provided in section E.5. First, we consider fine-tuning MACE-OFF (large) (Kovács et al., 2023), a MACE model pre-trained on the SPICE data set(Eastman et al., 2023), including QMugs and liquid water subsets (Isert et al., 2022; Schran et al., 2021). The foundation model is fine-tuned using the aspirin data set with energies and forces evaluated at the CCSD level of theory (Chmiela et al., 2018). We consider a more challenging scenario, e.g., where CC energies are extrapolated to the CBS limit. Thus, we use only energy labels for training.
We follow the fine-tuning procedure outlined in the official repository [^4] https://mace-docs.readthedocs.io/en/latest/guide/finetuning.html, adjusting the learning rate from $10^{-2}$ to $10^{-3}$ to enhance final accuracy of the model. The results, presented in Table 3, show that PIWSL significantly improves model accuracy, achieving approximately 40 % lower RMSE values. Notably, the model trained with PIWSL on 512 samples performs similarly to the model trained without it using nearly doubled training set size, i.e., 950 configurations. These results highlight the data efficiency of our approach. Furthermore, both fine-tuned MACE-OFF models outperform models trained from scratch.
Table 3: Results for models trained on the MD17(CCSD) data set without reference atomic forces. All models are trained on aspirin data without force labels. Energy RMSE is given in kcal/mol, while force RMSE is in kcal/mol/Å.
Model Samples Epoch Baseline PIWSL
MACE-OFF 950 100 E 1.21 $\pm$ 0.00 0.72 $\pm$ 0.01
Fa 6.90 $\pm$ 0.01 3.77 $\pm$ 0.13
MACE-OFF 512 100 E 2.03 $\pm$ 0.03 1.21 $\pm$ 0.02
Fa 10.96 $\pm$ 0.13 6.55 $\pm$ 0.27
MACE (scratch) 950 1000 E 3.27 $\pm$ 0.16 2.10 $\pm$ 0.30
Fa 18.05 $\pm$ 0.82 10.99 $\pm$ 0.18
a We evaluated the accuracy of predicted atomic forces using corresponding force labels provided in the test data set.
We also evaluate the impact of PIWSL on data sets containing conformations of a single large molecule. For this purpose, we have chosen the buckyball catcher molecule from the MD22 data set (Chmiela et al., 2023) with 148 atoms. We used two foundation models: MACE-MP (large) (Batatia et al., 2023) and MACE-OFF (large) (Kovács et al., 2023). We considered fine-tuning on a sparse data set; thus, both models were trained using only 50 samples. The results, presented in Table 4, demonstrate again that PIWSL improves the fine-tuning of foundation models. However, the performance of MACE-MP is worse compared to MACE-OFF, highlighting the importance of careful selection of the pre-trained model (Kjeldal & Eriksen, 2024). Note that a more rigorous hyper-parameter search could further enhance the performance of PIWSL. However, this is beyond the scope of this paper, as our primary objective is to demonstrate the relative effectiveness of the PIWSL approach.
Table 4: Energy (E) and force (F) mean-absolute errors (MAEs) for MACE, MACE-OFF, and MACE-MP models fine-tuned on the MD22 data set. All models are fine-tuned using 50 training samples of the buckyball-catcher molecule from MD22. Energy MAE is given in kcal/mol, while force MAE is in kcal/mol/Å.
Model Train Epoch Baseline PIWSL
MACE (scratch) 800 E 1.046 $\pm$ 0.095 0.745 $\pm$ 0.031
F 0.294 $\pm$ 0.002 0.290 $\pm$ 0.002
MACE-MP 100 E 2.111 $\pm$ 0.453 2.033 $\pm$ 0.149
F 0.716 $\pm$ 0.009 0.700 $\pm$ 0.011
MACE-OFF 100 E 1.159 $\pm$ 0.154 0.992 $\pm$ 0.051
F 0.346 $\pm$ 0.003 0.337 $\pm$ 0.001
@@ 6 Discussion and Limitations
This work introduces the PIWSL method, encompassing two distinct physics-informed weakly supervised loss functions, for learning MLIPs. These losses provide the physics-informed weak labels based on the Taylor expansion (PITC loss) and the spatial consistency (PISC loss) of the potential energy. These physics-informed weak labels enable any MLIP to improve its accuracy and robustness, particularly in scenarios characterized by sparse training data, which are common when investigating a new molecular or material system. The improved accuracy and robustness of MLIPs can allow running sufficiently long MD simulations, resulting in a more effective use of active learning approaches. Our extensive experiments demonstrate notable efficacy and efficiency of our method from various aspects: (i) dependence on the training data set size, (ii) the potential energy prediction variance and robustness in terms of a perturbation on a C–H bond length as well as robustness during MD simulations, and (iii) application to the fine-tuning of foundation models. In particular, it is shown that our PIWSL method enables MLIPs to improve the accuracy of their force predictions even without force labels used explicitly during training. Therefore, PIWSL opens a new possibility for training MLIPs using highly accurate reference methods, such as CCSD(T)/CBS.
Limitations. The proposed PIWSL method is tailored to ML models that predict atomic forces and total energies of atomic systems. It cannot be applied to other ML problems unrelated to computational chemistry or materials science. Although this work uses up to the second-order Taylor expansion to obtain weak labels in Eq. (4.1), employing more sophisticated higher-order ordinary differential equations is a viable alternative.
@@ Acknowledgements
Appendix A Potential Broader Impact and Ethical Aspects
Appendix B Related Work
Addressing Data Sparsity in MLIPs. Generating training data sets suitable for learning reliable MLIPs is challenging, especially when considering unexplored molecular and material systems. Numerous computationally expensive calculations with either ab initio or first-principles approaches are required for the latter. To mitigate this challenge, active learning (AL) methods, which utilize prediction uncertainty, can be applied (Li et al., 2015; Podryabinkin & Shapeev, 2017; Vandermause et al., 2020; Shuaibi et al., 2021b; Briganti & Lunghi, 2023; Zaverkin et al., 2024b). Furthermore, equivariant MLIPs often reduce required training data set sizes through improved data efficiency (Batzner et al., 2022; Batatia et al., 2022).
Appendix C Experimental Setup, Baselines, and Data Sets
C.1 Experimental Setup
Code for Experiments. The code used to run our experiments builds upon the recent work (Fu et al., 2023) and extends it to integrate the latest Open Catalyst Project code (Chanussot et al., 2021). We adopt hyper-parameters from the Open Catalyst (OC) project, tuned to the corresponding OC data set. Note that we do not use this data set in the presented work, whose main focus is training general-purpose MLIPs that can be used to run molecular dynamics (MD) simulations and geometry optimization. However, the OC data set has been designed to investigate the latter, making it less suitable for the current study. Our modifications include adjusting the learning rate scheduler, details of which can be found in our repository. For potential energy and force prediction, we utilize mean-absolute error (MAE) and $L_{2}$-norm (L2MAE) losses with coefficients of 1 and 100, respectively. More details on the model hyperparameters are provided in our repository. For the PITC and PISC loss functions, we use the mean square error (MSE) loss based on an experiment in section E.6.
@@ C.1 Experimental Setup
a As explained in Table 2, the mini-batch size of the SchNet model was changed to 32 due to high RMSE values observed with a mini-batch size of four as the training data set size increased. A more detailed discussion of the results for SchNet is provided in section E.1.
@@
The remaining hyper-parameters are the coefficients for the PITC and PISC losses ($C_{\rm PITC},C_{\rm PISC}$) and the maximum magnitude $\epsilon_{\mathrm{max}}$ of the perturbation vector $\delta\mathbf{r}$; see Table A3. These hyper-parameters are tuned using Optuna (Akiba et al., 2019) for PaiNN and Equiformer v2. We used 1000 configurations drawn randomly from the original ANI-1x data set for training. Due to multiple local minima, Optuna identified several optimal hyper-parameter sets in each run. We selected the following representative combinations ($C_{\rm PITC},C_{\rm PISC},\epsilon_{\mathrm{max}})$ = Case A: (1.2, 0.8, 0.025), Case B: (1.0, 0, 0.01), Case C: (0.1, 0.01, 0.01), Case D: (1.2, 0.01, 0.025), Case E: (1.2, 0.01, 0.01), Case F: (1.2, 0.01, 0.015), Case G: (0.01, 0.001, 0.025), and Case H: (0.1, 0.01, 0.025). We selected the hyper-parameters listed in Table A3 based on the validation dataset performance.
Table A3: Hyper-parameters for the PIWSL loss. We selected the following hyper-parameter combinations using Optuna (Akiba et al., 2019): ($C_{\rm PITC},C_{\rm PISC},\epsilon_{\mathrm{max}})$= Case A: (1.2, 0.8, 0.025), Case B: (1.0, 0, 0.01), Case C: (0.1, 0.01, 0.01), Case D: (1.2, 0.01, 0.025), Case E: (1.2, 0.01, 0.01), Case F: (1.2, 0.01, 0.015), Case G: (0.01, 0.001, 0.025), and Case H: (0.1, 0.01, 0.025).
@@
Splitting Data Sets. We split the original data sets into training, validation, and test sets for our experiments. We shuffled the original data sets using a random seed and selected the training data sets of predefined sizes. For validation, we selected the same number of configurations as in the training data set if it exceeded 100 configurations; otherwise, we used 100 configurations to ensure sufficient validation size. For the rMD17 data set, following (Fu et al., 2023), we used 9000 configurations as a validation data set and another 10,000 for testing. We used the same test data set across different sizes of the training data sets for a fair performance comparison. We used 10,000 test configurations for ANI-1x and 1000 for TiO2 and LMNTO.
Training with Adversarial Directions. In our experiments, which defined perturbation vectors adversarially (see section F), we determined adversarial directions using Eq. (7). More concretely, we only considered the potential energy, i.e., $\mathbf{y}_{\rm pred}$ and $\mathbf{y}_{\rm label}$, to avoid Hessian calculations. In addition, we considered the loss function for the potential energy as $L_{\rm dist}$ in Eq. (7). The expression of $\mathbf{g}=\nabla_{\mathbf{r}}L_{\rm dist}$ is then
@@
Note that we used the relation $\nabla_{\mathbf{r}_{i}}E^{\rm ref}_{\mathcal{S}}=-\mathbf{F}^{\rm ref}_{i,\mathcal{S}}$ to obtain the final expression. Though $E^{\rm ref}$ can also be interpreted as a constant regarding atom positions. We have chosen this expression to avoid the case where the adversarial direction $\mathbf{g}$ points to $\mathbf{F}^{\rm ref}_{i,\mathcal{S}}$ other than the very beginning of the training. Our experiments indicate that the employed expression is slightly better than its alternative. In our experiment, we also randomly flip the sign of $\mathbf{g}_{\mathcal{S}}$ to avoid overfitting to adversarial directions.
C.2 Baseline Methods
Data Augmentation with NoisyNode. In our experiments, we used the NoisyNode approach (Godwin et al., 2022) as one of the baseline methods. This method aims to improve the performance of ML models by adding a perturbation to node features, i.e., atomic coordinates, and makes ML models recover original labels. This approach enables ML models to be more robust to noise in the data. Although the original method recommends adding a decoder network to learn the denoising process, we do not utilize it following previous work (Liao et al., 2023) and add the perturbation vector to atomic coordinates similar to PIWSL losses, fixing energy and force labels. We implement the NoisyNode approach in our code. Thus, we can expect slightly different behavior compared to the recent work (Godwin et al., 2022; Liao et al., 2023)  [^5] Note that NoisyNode assumes the unperturbed state to be the equilibrium structure, which may have contributed to the limited performance improvements observed in our experiments. Recently, efforts have been underway to develop an extended version of NoisyNode tailored for non-equilibrium structures, such as (Liao et al., 2024)..
Taylor-Expansion-Based Weak Labels. Recent work proposed a similar Taylor-expansion-based weak label approach (Cooper et al., 2020). Nonetheless, the loss is different from the one in Eq. (4.1) as the authors used reference energy and atomic force labels to estimate weak energy labels $E^{\mathrm{ref}}_{\mathcal{S}_{\delta\mathbf{r}}}$ for perturbed atomic configurations $\mathcal{S}_{\delta\mathbf{r}}$